# Categories of Line Defects and Cohomological Hall Algebras

### v1

@@ Abstract
Any four-dimensional Supersymmetric Quantum Field Theory with eight supercharges can be associated to a monoidal category of BPS line defects. Any Coulomb vacuum of such a theory can be conjecturally associated to an “algebra of BPS particles”, exemplified by certain Cohomological Hall Algebras. We conjecture the existence of a monoidal functor from the category of line defects to a certain category of bimodules for the BPS Algebra in any Coulomb vacuum. We describe images of simple objects under the conjectural functor and study their monoidal structure in examples. As we vary the choice of vacuum, we expect the collection of functors associated to any given theory to capture the full information of the original monoidal category of lines.
@@ 1 Introduction
Four-dimensional ${\cal N}=2$ Supersymmetric Quantum Field Theories typically admit Coulomb vacua whose massless spectrum consists of a free supersymmetric Abelian gauge theory Seiberg:1994rs ; Seiberg:1994aj ; Klemm:1994qs ; Argyres:1994xh ; Hanany:1995na ; Argyres:1995jj ; Martinec:1995by ; Argyres:1996eh ; Katz:1996fh ; Witten:1997sc ; Gaiotto:2009we . The Coulomb vacua form a complex manifold ${\mathcal{B}}$, with singular loci where extra light degrees of freedom emerge. The gauge couplings of the IR gauge fields vary holomorphically on ${\mathcal{B}}$, up to electric-magnetic dualities. The non-perturbative physical insights encoded in the structure of the Coulomb vacua are crucial for many mathematical applications of 4D ${\cal N}=2$ SQFTs Witten:1994cg ; Donagi:1997sr ; Nekrasov:2002qd ; Alday:2009aq ; Nekrasov:2009rc ; Gaiotto:2010okc .
@@
In the physics literature, the BPS spectrum was initially known only in a few examples, collating information about singular loci in ${\mathcal{B}}$ where BPS particles become massless and the free Abelian low energy description is incomplete. The Kontsevich-Soibelman wall-crossing formula Kontsevich:2008fj greatly improved our understanding by providing a precise recipe for the $u$ dependence of the BPS spectrum.
@@
There are strong indications Li:2023zub that framed quivers may provide a link between the IR behaviour of half-BPS line defects and the representation theory of CoHA’s. The main objective of this paper is to propose and sharpen the following conjecture:
@@ 1.1 Half-BPS lines and the Holomorphic-Topological twist
  • Upon compactification on a circle, wrapped half-BPS lines map to linear combinations of wrapped Abelian ’t Hooft-Wilson lines. This map endows ${\mathcal{M}}$ with the structure of a cluster variety.
@@
This mathematical formalization of UV physics does not include information about the RG flow or the notion of a Coulomb branch vacuum. This is similar to what happens to the B-twist of 2D $(2,2)$ SQFTs: the category of D-branes is known to coincide with the derived category of coherent sheaves, but the effect of RG flows on the category of branes is poorly understood and challenging to compute mathematically unless a mirror dual is available.
@@
Accordingly, the image $\mathrm{RG}_{u}(L)$ is essentially given by the Chan-Paton coefficients [^2] To be precise, the Chan-Paton coefficients should carry a representation of some kind of Yangian algebra for the Abelian HT theory Costello:2013zra , Koszul dual to the fermionic ghosts. We briefly discuss a related symmetry in Section 2.8 but do not attempt a systematic treatment. $V_{\gamma;n}[L]$ in
@@
We are inspired by the setting of gapped 2D TFTs, where the “Algebra of the Infrared” formalism Gaiotto:2015aoa builds IR categories that are conjecturally isomorphic to the categories of boundary conditions or interfaces of the microscopic theory. BPS particles play a crucial role in the 2D story and we expect them to play a crucial role here as well.
@@
is actually the object-level part of a monoidal functor into the category of $A_{u}$-bimodules.
We present two pieces of evidence for our proposal:
@@
    We verify in simple non-Lagrangian examples that the tensor product of $B_{L;u}$ bimodules indeed can be written as an extension of $B_{L;u}$’s, and that the structure of the extensions is compatible with the existence of functors from an $u$-independent category.
@@ 1.2 Future directions and missing links
A traditional way to “improve” a fiber functor is to include information about the action of the algebra of endomorphisms of the fiber functor itself private . This is the idea of Koszul duality Costello:2016mgj , which was used to great effect in other twisted theories Costello:2013zra ; Paquette:2021cij . Despite tantalizing hints that the algebra of endomorphisms should be a “double” Rapcak:2018nsl of the CoHA $A_{u}$, we haven’t been able to bring this strategy to fruition yet.
Another related, incomplete circle of ideas that was an important motivation for this work concerns the “Schur index” of the 4D theory and its IR presentation. The Schur index can be understood as a graded Witten index for the space of local operators of the HT twisted theory. The Schur index can be modified by line operator insertions to compute what is essentially a graded Witten index for the space of morphisms between two line defects.
The Schur index of any theory which admits a Coulomb vacuum can be reconstructed from the spectrum of BPS particles in that vacuum and from the framed BPS indices of the line defects Cordova:2015nma ; Cordova:2016uwk . The formula has a form familiar from 2D: it appears to combine the naive morphisms in $\mathrm{Line}^{IR}_{u}$ with a count of “fans” of BPS particles emanating from a point with slopes controlled by the phases of their “central charges”.
Accordingly, one may hope to build an “IR description” of $\mathrm{Line}$ by formalizing the notion of fans of BPS particles and a way to compose them. A previous work Gaiotto:2023dvs suggested that spaces of “half fans” comprised of BPS particles emanating out in an half-plane could be recovered from $A_{u}$ via Koszul duality. Again, a consistent procedure to double the space of half-fans to a space of fans eluded us.
In 2D, half-fans give the endomorphism algebra of a generating object in the category of boundary conditions. Interfaces can be described alternatively as bimodules for the algebra of half-fans or its Koszul dual, if available. This somewhat contrived path may lead one to study $A_{u}$-bimodules, as we do. We hope that the correct formulation of the category of CoHA bimodules we study in this paper will make contact with IR formulae for the Schur index.
@@ 2.3 Stability conditions and walls
The stability conditions are controlled by the central charges $Z_{\gamma}(u)$, which define a linear map $\Gamma_{u}\to\mathbb{C}$. The central charges are locally holomorphic and non-vanishing on the non-singular part of ${\mathcal{B}}$. In the rest of the paper we will never really need to refer directly to $u$: only $\Gamma$ and $Z_{\gamma}$ enter our considerations. We will thus drop any reference to $u$ in the following.777 In general, not all values of the central charges occur in ${\mathcal{B}}$. Accordingly, not all possible quiver mutations will occur along ${\mathcal{B}}$ Cecotti:2011rv . The implications of this fact are unclear.
@@
for both $\mu^{0/\pi}_{i}$. For the same $i$, the two mutations $\mu^{0}_{i}$ and $\mu^{\pi}_{i}$ are inverses of each other, and the square of a mutation only changes the labelling of a quiver, by the Lefshetz monodromy:
@@ 2.4 Heisenberg grading and Kontsevich-Soibelman spectrum generator
for the tensor product. Equivalently,
$(V\otimes W)_{\gamma,r}=\oplus_{\gamma^{\prime}\in\Gamma,r^{\prime}\in\mathbb{Z}}V_{\gamma^{\prime},r^{\prime}}\otimes W_{\gamma-\gamma^{\prime},r-r^{\prime}+\langle\gamma,\gamma^{\prime}\rangle}\,.$ (21)
@@ 2.6 Spherical CoHA and Spherical Shuffle Algebra
Physically, each node of the quiver represents a type of D-branes/BPS particle of charge $\gamma_{i}$, and the $s_{i,a}$ variables can be reasonably interpreted as the positions of $n_{i}$ indistinguishable particles in the $\mathbb{C}$ direction of the HT twist. The polynomial is an holomorphic analogue of a wavefunction. See Appendix B for details.
@@ 2.8 Extra symmetries
We can denote the combination of the usual Heisemberg grading and these extra symmetry transformations as $\Gamma_{\mathrm{Hei}}^{\vee}({\mathcal{O}})$. [^13] If the $\gamma_{i}$ do not generate $\Gamma$, we are adding some extra symmetry generators which act trivially on the algebra.
@@ 3.2 The $A_{3}$ quiver and its relatives
    We do not know how to prove directly that this is indeed the character of $\mathrm{sH}(\hat{A}_{3})$, but we have computed by hand the dimensions of the graded components $\mathrm{sH}_{\gamma,k}(\hat{A}_{2})$ for small values of the charges and found agreement with the above expression. We have also identified two candidates for the fermionic generators in $S_{\gamma_{2}+\gamma_{3}}$:
@@ 4.5 Tensor products and factorization
We can thus refine our conjecture. We expect that it should be possible to define a monoidal category $\mathrm{sH}(Q)-\mathrm{Bimod}^{\mathrm{PBW}}$ of “PBW-bimodules”, which admit compatible PBW factorizations for each $\vartheta$ and only include morphisms which are induced from morphisms of the $V$’s.
@@
The bimodules that we have defined in this section also enjoy the obvious action of the $\Gamma_{\mathrm{Hei}}^{\vee}({\mathcal{O}})$ symmetry algebra, which commutes with the bimodule action. Our tensor product decompositions are also compatible with the $\Gamma_{\mathrm{Hei}}^{\vee}({\mathcal{O}})$ action.
@@
    $\mathrm{RG}_{u}:\qquad\mathrm{Lines}\,\to\,\mathrm{sH}(Q)-\mathrm{Bimod}^{\mathrm{PBW}}_{\Gamma_{\mathrm{Hei}}^{\vee}({\mathcal{O}})}\,,$ (154)
    mapping simple objects $L$ to $\mathrm{sH}(Q;\langle\gamma_{*}(L),\vec{\gamma}\rangle)$, and involving a category of bimodules that is both $\Gamma_{\mathrm{Hei}}^{\vee}({\mathcal{O}})$-equivariant and compatible with PBW factorizations.

### v2

@@ Abstract
Any four-dimensional Supersymmetric Quantum Field Theory with eight supercharges can be associated to a monoidal category of BPS line defects. Any Coulomb vacuum of such a theory can be conjecturally associated to an “algebra of BPS particles”, exemplified by certain Cohomological Hall Algebras. We conjecture the existence of a monoidal functor from the category of line defects to a certain category of bimodules for the BPS Algebra in any Coulomb vacuum. We describe images of simple objects under the conjectural functor and study their monoidal structure in examples. We conjecture that the functor may be an equivalence of dg-categories and test the conjecture at the level of the equivariant Witten indices of the spaces of morphisms.
@@ 1 Introduction
Four-dimensional ${\cal N}=2$ Supersymmetric Quantum Field Theories typically admit Coulomb vacua whose massless spectrum consists of a free supersymmetric Abelian gauge theory Seiberg:1994rs ; Seiberg:1994aj ; Klemm:1994qs ; Argyres:1994xh ; Hanany:1995na ; Argyres:1995jj ; Martinec:1995by ; Itoyama:1995nv ; Argyres:1996eh ; Katz:1996fh ; Witten:1997sc ; Gaiotto:2009we . The Coulomb vacua form a complex manifold ${\mathcal{B}}$, with singular loci where extra light degrees of freedom emerge. The gauge couplings of the IR gauge fields vary holomorphically on ${\mathcal{B}}$, up to electric-magnetic dualities. The non-perturbative physical insights encoded in the structure of the Coulomb vacua are crucial for many mathematical applications of 4D ${\cal N}=2$ SQFTs Witten:1994cg ; Donagi:1997sr ; Nekrasov:2002qd ; Alday:2009aq ; Nekrasov:2009rc ; Gaiotto:2010okc .
@@
In the physics literature, the BPS spectrum was initially known only in a few examples, collating information about singular loci in ${\mathcal{B}}$ where BPS particles become massless and the free Abelian low energy description is incomplete. The Kontsevich-Soibelman wall-crossing formula Kontsevich:2008fj greatly improved our understanding by providing a precise recipe for the $u$ dependence of the BPS spectrum. [^1] See also Denef:2007vg for an earlier physics story.
@@
There are strong indications Li:2023zub that framed quivers may provide a link between the IR behaviour of half-BPS line defects and the representation theory of CoHA’s. The main objective of this paper is to propose and sharpen the following conjectures:
@@
  • When a quiver description is available, the bimodules are built from framed quivers by a certain variant of the CoHA construction.
  • When properly defined, the functor is a quasi-isomorphism. In particular the categories associated to different vacua are also quasi-isomorphic to each other.
@@ 1.1 Half-BPS lines and the Holomorphic-Topological twist
  • Upon compactification on a circle, wrapped half-BPS lines map under RG flow to linear combinations of wrapped Abelian ’t Hooft-Wilson lines. This map often endows ${\mathcal{M}}$ with the structure of a cluster variety.
@@
Finally, the equivariant characters of the dg morphism spaces of the $\mathrm{Line}$ category coincide with the so-called Schur indices Gadde:2011uv ; Dimofte:2011py ; Cordova:2016uwk of the underlying 4d SQFT Niu:2021jet .
This mathematical formalization of UV physics does not include information about the RG flow or the notion of a Coulomb branch vacuum. This is similar to what happens to the B-twist of 2D $(2,2)$ SQFTs: the category of D-branes is known to coincide with the derived category of coherent sheaves, but the effect of RG flows on the category of branes is poorly understood and challenging to compute mathematically unless a mirror dual is available Herbst:2008jq ; Hori:2013ika .
@@
Accordingly, the image $\mathrm{RG}_{u}(L)$ is essentially given by the Chan-Paton coefficients [^3] To be precise, the Chan-Paton coefficients should carry a representation of some kind of Yangian algebra for the Abelian HT theory Costello:2013zra , Koszul dual to the fermionic ghosts. We briefly discuss a related symmetry in Section 2.8 but do not attempt a systematic treatment in that language. $V_{\gamma;n}[L]$ in
@@
We are inspired by the setting of gapped 2D TFTs, where the “Algebra of the Infrared” formalism Gaiotto:2015aoa builds IR categories that are conjecturally quasi-isomorphic to the categories of boundary conditions or interfaces of the microscopic theory. BPS particles play a crucial role in the 2D story and we expect them to play a crucial role here as well.
@@
is actually the object-level part of a monoidal functor into a category of $A_{u}$-bimodules which are equivariant under a certain gauge action.
We present three pieces of evidence for our proposal:
@@
    We verify in simple non-Lagrangian examples that the tensor product of $B_{L;u}$ bimodules indeed can be written as an extension of $B_{L;u}$’s, and that the structure of the extensions is compatible with the existence of functors from an $u$-independent ${\mathrm{Line}}$ category.
  3. 3.
    We verify that an appropriately-defined space of morphisms of bimodules has an $u$-independent graded character which coincides with the expected answer for the $\mathrm{Line}$ category (the Schur indices), compatibly with these functors being quasi-isomorphisms.
@@
The formula we derive for the graded character of morphism spaces use the proposal that the Schur indices of any theory which admits a Coulomb vacuum can be reconstructed from the spectrum of BPS particles in that vacuum and from the framed BPS indices of the line defects Cordova:2015nma ; Cordova:2016uwk .
@@ 1.2 Future directions and missing links
At the beginning of this project, our objective was to promote the $V_{L,u}$ spaces to modules for some algebra equipped with a co-product, perhaps a “double” Rapcak:2018nsl of the algebra of BPS states. Indeed, a traditional way to “improve” a fiber functor is to include information about the action of the algebra of endomorphisms of the fiber functor itself private . This is the idea of Koszul duality Costello:2016mgj , which was used to great effect in other twisted theories Costello:2013zra ; Paquette:2021cij .
We could not make this idea work and found instead the alternative route of a bimodule category. It would nevertheless be interesting to identify such a description for the category of line defects.
It would be also useful to find a clear physical explanation for our proposal. In 2d analogues, the category of line defects can be built directly from the IR data by assembling BPS particles into “fans” which can be created by local operators Gaiotto:2015aoa . A previous work Gaiotto:2023dvs suggested that spaces of “half fans” comprised of BPS particles emanating out in an half-plane could be recovered from $A_{u}$ via Koszul duality. We are not sure how to assemble these ideas into a coherent whole.
@@ 2.3 Stability conditions and walls
The stability conditions are controlled by the central charges $Z_{\gamma}(u)$, which define a linear map $\Gamma_{u}\to\mathbb{C}$. The central charges are locally holomorphic and non-vanishing on the non-singular part of ${\mathcal{B}}$. In the rest of the paper we will never really need to refer directly to $u$: only $\Gamma$ and $Z_{\gamma}$ enter our considerations. We will thus drop any reference to $u$ in the following.888 In general, not all values of the central charges occur in ${\mathcal{B}}$. Accordingly, not all possible quiver mutations will occur along ${\mathcal{B}}$ Cecotti:2011rv . The implications of this fact are unclear.
@@
for both $\mu^{0/\pi}_{i}$. For the same $i$, the two mutations $\mu^{0}_{i}$ and $\mu^{\pi}_{i}$ are inverses of each other, and the square of a mutation only changes the labelling of a quiver, by the Lefschetz monodromy:
@@ 2.4 Heisenberg grading and Kontsevich-Soibelman spectrum generator
for the tensor product. [^10] Physically, this is the contribution of the IR electric-magnetic fields to the spin of a system of parallel dyons. Equivalently,
$(V\otimes V^{\prime})_{\gamma,r}=\oplus_{\gamma^{\prime}\in\Gamma,r^{\prime}\in\mathbb{Z}}V_{\gamma^{\prime},r^{\prime}}\otimes V^{\prime}_{\gamma-\gamma^{\prime},r-r^{\prime}+\langle\gamma,\gamma^{\prime}\rangle}\,.$ (21)
@@ 2.6 Spherical CoHA and Spherical Shuffle Algebra
Physically, each node of the quiver represents a type of D-branes/BPS particle of charge $\gamma_{i}$, and the $s_{i,a}$ variables can be reasonably interpreted as the positions of $n_{i}$ indistinguishable particles in the $\mathbb{C}$ direction of the HT twist. The polynomial is a holomorphic analogue of a wavefunction. See Appendix B for details.
@@ 2.8 Extra symmetries
We can denote these extra symmetry transformations as $\Gamma^{\vee}({\mathcal{O}})$ and the combination with the usual Heisenberg grading as $\Gamma_{\mathrm{Hei}}^{\vee}({\mathcal{O}})$. [^15] If the $\gamma_{i}$ do not generate $\Gamma$, we are adding some extra symmetry generators which act trivially on the algebra.
@@ 3.2 The $A_{3}$ quiver and its relatives
    We do not know how to prove directly that this is indeed the character of $\mathrm{sH}(\hat{A}_{2})$, but we have computed by hand the dimensions of the graded components $\mathrm{sH}_{\gamma,k}(\hat{A}_{2})$ for small values of the charges and found agreement with the above expression. We have also identified two candidates for the fermionic generators in $S_{\gamma_{2}+\gamma_{3}}$:
@@ 4.5 Tensor products and factorization
We can thus refine our conjecture. We expect that it should be possible to define a monoidal category $\mathrm{sH}(Q)-\mathrm{Bimod}^{\mathrm{PBW}}$ of “PBW-bimodules”, which admit compatible PBW factorizations for each $\vartheta$. It seems to be some kind of dualizability condition on the bimodules.
@@
4.6 IR gauge equivariance
The bimodules that we have defined in this section also enjoy the obvious action of the $\Gamma^{\vee}({\mathcal{O}})$ symmetry algebra, which commutes with the bimodule action. Our tensor product decompositions are also compatible with the $\Gamma^{\vee}({\mathcal{O}})$ action.
@@ 4.6 IR gauge equivariance
    $\mathrm{RG}_{u}:\qquad\mathrm{Lines}\,\to\,\mathrm{sH}(Q)-\mathrm{Bimod}^{\mathrm{PBW}}_{\Gamma^{\vee}({\mathcal{O}})}\,,$ (154)
    mapping simple objects $L$ to $\mathrm{sH}(Q;\langle\gamma_{*}(L),\vec{\gamma}\rangle)$, and involving a category of bimodules that is both $\Gamma^{\vee}({\mathcal{O}})$-equivariant and compatible with PBW factorizations.
This final refinement will turn out to be crucial to reproduce the desired categorical structures.
4.7 A test of invertibility for $\mathrm{RG}_{u}$
We will now compute the graded Witten index of the $\mathrm{Hom}$ spaces in a category of $\Gamma^{\vee}({\mathcal{O}})$-equivariant bimodules.
If we ignore momentarily the $\Gamma^{\vee}({\mathcal{O}})$ equivariance, the Witten index can be computed from a bar complex for the derived spaces of morphisms. Schematically, the complexes are the direct sum of spaces of multi-linear maps
$(A/\mathbb{C}1)^{\otimes n}\otimes B_{1}\otimes(A/\mathbb{C}1)^{\otimes m}\to B_{2}$ (155)
where $B_{1}$ and $B_{2}$ are bimodules, $(A/\mathbb{C}1)$ is the algebra with the identity quotiented away and the summand is shifted in ghost number by $n+m$. The character of either of the infinite direct sums gives
$\frac{1}{1+(\chi(A)-1)}=\chi(A)^{-1}\,.$ (156)
and effectively “cancels” a copy of $A$.
The overall structure of the calculation is already visible from the computation of $\mathrm{End}(A)$:
$\chi(\mathrm{End}_{A-\mathrm{Bimod}}(A))=\chi(A)^{-1}\chi(A)\chi(A)^{-1}\chi(A^{\vee})=\chi(A)^{-1}\chi(A^{\vee})$ (157)
We impose $\Gamma^{\vee}({\mathcal{O}})$-equivariance in a relative sense: we impose $\Gamma^{\vee}$ invariance by hand by projecting on zero total gauge charge and add $c_{i,n}$ ghosts for the remaining generators:
$\chi(\mathrm{End}_{A-\mathrm{Bimod}_{\Gamma^{\vee}({\mathcal{O}})}}(A))=(\mathfrak{q}^{2})^{\mathrm{rk}(\Gamma)}_{\infty}\mathrm{Tr}[\chi(A)^{-1}\chi(A^{\vee})]\,,$ (158)
where “$\mathrm{Tr}$” denotes the projection $X_{\gamma}\to\delta_{\gamma,0}$.
Recall that $\chi(A)={\mathcal{S}}(\mathfrak{q})$. Analogously, $\chi(A^{\vee})={\mathcal{S}}(\mathfrak{q}^{-1})^{t}$, where the $t$ superscript denotes $X_{\gamma}\to X_{-\gamma}$. We now conjecture a marvellous relation:
${\mathcal{S}}(\mathfrak{q}^{-1})={\mathcal{S}}^{-1}(\mathfrak{q})\,.$ (159)
This identity holds automatically when ${\mathcal{S}}$ is a product of quantum dilogarithms with quantum torus arguments, as it holds for the individual factors:
$\Phi(x;\mathfrak{q}^{-1})=\Phi(x;\mathfrak{q})^{-1}\,,$ (160)
and both inversion and $\mathfrak{q}\to\mathfrak{q}^{-1}$ invert the order of multiplication in the quantum torus algebra.
Then
$\chi(\mathrm{End}_{A-\mathrm{Bimod}_{\Gamma^{\vee}({\mathcal{O}})}}(A))=(\mathfrak{q}^{2})^{\mathrm{rk}(\Gamma)}_{\infty}\mathrm{Tr}[{\mathcal{S}}^{-1}({\mathcal{S}}^{-1})^{t}]\,.$ (161)
This is the IR formula for the Schur index Cordova:2015nma ; Cordova:2016uwk !
The Schur index computes the Witten index of the space of local operators of the UV theory, i.e.
$\chi(\mathrm{End}_{A-\mathrm{Bimod}_{\Gamma^{\vee}({\mathcal{O}})}}(A))=\chi(\mathrm{End}_{\mathrm{Lines}}(1))\,.$ (162)
We can do better:
$\chi(\mathrm{Hom}_{A-\mathrm{Bimod}}(B_{1},B_{2}))=\chi(A)^{-1}\chi(B_{2})\chi(A)^{-1}\chi(B_{1}^{\vee})=\chi(A)^{-1}\chi(V_{2})\chi(A^{\vee})\chi(V_{1}^{\vee})$ (163)
and thus for bimodules $B_{L_{i}}$ which are $\mathrm{RG}_{u}$ images of $L_{i}$ we have
$\chi(\mathrm{Hom}_{A-\mathrm{Bimod}_{\Gamma^{\vee}({\mathcal{O}})}}(B_{L_{1}},B_{L_{2}}))=(\mathfrak{q}^{2})^{\mathrm{rk}(\Gamma)}_{\infty}\mathrm{Tr}[({\mathcal{S}}^{-1}F^{\pi}_{L_{2}})({\mathcal{S}}^{-1}F^{\pi}_{L_{1}})^{t}]\,,$ (164)
where we have assumed the Poincaré duality that the $F_{L}$ are symmetric under $\mathfrak{q}\to\mathfrak{q}^{-1}$. This is the IR formula for the Schur index counting defect-changing local operators Cordova:2015nma ; Cordova:2016uwk , i.e.
$\chi(\mathrm{Hom}_{A-\mathrm{Bimod}_{\Gamma^{\vee}({\mathcal{O}})}}(B_{L_{1}},B_{L_{2}}))=\chi(\mathrm{Hom}_{\mathrm{Lines}}(L_{1},L_{2}))=\chi(\mathrm{Hom}_{\mathrm{Lines}}(1,\rho(L_{1})L_{2}))\,,$ (165)
where in the last step we have used the automorphism $\rho$ that rotates the lines by $\pi$, which satisfies $F^{\pi}_{\rho(L)}=(F^{0}_{L})^{t}$ and $\textrm{Tr}[ab]=\textrm{Tr}[\rho^{2}(b)a]$.
These formulae are fully consistent with the hypothesis that $\mathrm{RG}_{u}$ is a quasi-isomorphism at the level of $\mathrm{Hom}$ spaces!
4.8 Further constraints
At this point, we have a natural IR image for each $L_{i}$ and a plausible expectation of a quasi-isomorphism for the corresponding morphism spaces.
It would be nice to find a stringent-enough set of constraints which could restrict the bimodule category so that the $\mathrm{RG}_{u}$ functors are surjective as well.
We have already observed the existence of a PBW factorization for the $B_{L_{i}}$ and their tensor products. We should mention another property which appears to be common to all of these bimodules.
Observe that the exchange relations allow one to bring generators $e_{n}^{(i)}$ across any element in the bimodules as long as $n$ is sufficiently large, up to a shift of the $n$ argument. One may include such property as a further constraint on the IR bimodule category.